OpenStudy (anonymous):

-2x-y=9 x+7y=15 Substitution method I got x= 34/13?

OpenStudy (anonymous):

from equation 2: x=15-7y subbing into equation 1: -30+14y-y=9 y=39/13=3 subbing back into equation 2: x+21=15 x=-6

OpenStudy (dan815):